MICROWAVE SPECTRA AND STRUCTURES OF H_{2}O-AgCl AND H_{3}N-AgCl.

\underline{N. R. WALKER}, V. A. MIKHAILOV, S. J. HARRIS, F. J. ROBERTS AND A. C. LEGON, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K..

A Balle-Flygare FT-MW spectrometer coupled to a laser ablation source has been used to measure the pure rotational spectra of H_{2}O-AgCl and H_{3}N-AgCl. Both molecules are generated via laser ablation (532 nm) of a metal rod in the presence of CCl_{4}, argon, a low partial pressure of H_{2}O or NH_{3} and are stabilized by supersonic expansion. Rotational constants and centrifugal distortion constants have been measured for eight isotopologues of H_{2}O-AgCl with substitutions at the silver, chlorine, oxygen and hydrogen atoms. B_{0} and C_{0} have been independently determined for six isotopologues and the spectra are consistent with a linear arrangement of oxygen, silver and chlorine atoms. The structure is either C_{2v} planar at equilibrium or C_{S} pyramidal but with a low potential-energy barrier to planarity such that the v=0 and 1 states associated with the motion that inverts the configuration at the O atom are well separated. Preliminary data from a study of H_{3}N-AgCl will also be presented. Several transitions have been observed and tentatively assigned to hyperfine components of the 2_{02}-1_{01}, 3_{03}-2_{02} and 4_{04}-3_{03} transitions of H_{3}N-^{107}Ag^{35}Cl and H_{3}N-^{109}Ag^{35}Cl. These data are consistent with a C_{3v} structure.